OpenStudy (anonymous):

find general solution to y prime +(4/x-1)y = 1/(x-1)^5 + sin(x)/(x-1)^4

OpenStudy (anonymous):

\[y \prime + \frac{ 4 }{ (x-1) }y = \frac{ 1 }{ (x-1)^{5} }+\frac{ \sin(x) }{ (x-1)^{4} }\]

OpenStudy (anonymous):

I'm totally lost and I don't know how to even go about this

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

ughhhhhhhhh math

ganeshie8 (ganeshie8):

multiply throught out by (x-1)^4 and see something magic happens on left hand side

OpenStudy (anonymous):

\[(x-1)^{4} y \prime + f(x-1)^{3} y\] ?

OpenStudy (anonymous):

does y prime cancel out?

OpenStudy (anonymous):

i have no idea, but wolfram does looks like @ganeshie8 had the right idea

OpenStudy (anonymous):

I still don't understand the procedure :(

ganeshie8 (ganeshie8):

the trick is to multiply something both sides so that the left hand gets into form : \[\large (y*g)'\]

ganeshie8 (ganeshie8):

after that, integrating both sides simplifies the left hand side to \(\large y*g\) because the integral and derivative eat eachother out

ganeshie8 (ganeshie8):

that special factor to be multiplied both sides hapens to be (x+1)^4 here

ganeshie8 (ganeshie8):

\[\large y' + \frac{ 4 }{ (x-1) }y = \frac{ 1 }{ (x-1)^{5} }+\frac{ \sin(x) }{ (x-1)^{4} } \] Multiply (x-1)^4 both sides : (there is a standard way to find this factor) \[ \large (x-1)^4y' + 4(x-1)^3y = \frac{ 1 }{ x-1 }+\sin x \]

ganeshie8 (ganeshie8):

stare at left hand side, does it look familiar : \( \large fg' + f'g\)

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

@ganeshie8 i have a question was this cooked up so this gimmick works for this question, or is there something in the method that forces it?

ganeshie8 (ganeshie8):

yes use it to your advantage here, write left hand side as (fg)'

ganeshie8 (ganeshie8):

satellite, it works in general for equations of form : \[\large y' + f(x) y = g(x)\]

ganeshie8 (ganeshie8):

we can always find such a factor, multiply it both sides, and write the left hand side as derivative of product of two functions : (yg)'

OpenStudy (anonymous):

and then integrate?

ganeshie8 (ganeshie8):

yes integrating isolates the required function "y" thats ur goal function when solving a differential equation

ganeshie8 (ganeshie8):

\[\large (x-1)^4y' + 4(x-1)^3y = \frac{ 1 }{ x-1 }+\sin x \] rewriting left hand side using reverse product rule you get \[\large \left((x-1)^4y\right)' = \frac{ 1 }{ x-1 }+\sin x \]

ganeshie8 (ganeshie8):

integrate both sides and solve y

OpenStudy (anonymous):

okay thanks I'm going to try to solve it

ganeshie8 (ganeshie8):

rest of the problem is straightforward : \[\large \int \left((x-1)^4y\right)' dy= \int \frac{ 1 }{ x-1 }+\sin x dx \] \[\large (x-1)^4y= \int \frac{ 1 }{ x-1 }+\sin x dx \] evaluate the right hand side integral and you're done!

OpenStudy (anonymous):

So just to verify, \[y=\frac{ -1 }{ 4 }-\frac{ \cos(x) }{ (x-1)^4 }\]